# Renormalon-like factorial enhancements to power expansion/OPE expansion in super-renormalizable QFTs

### v1

@@ 4 Comments
    First, we emphasize that due to the off-diagonal cancellation pattern across different powers, the naive power-expansion in $\frac{1}{(p^{2})^{n+2}}$ for the two point function, although free from any ambiguity at any given power, failed to converge. The OPE expansion is also not convergent, unless one redefine all the operators by subtracting out all the factorial growth and add back to coefficient functions. Even after this procedure, the coefficient function still can not be summed as a power expansion.
@@
    We should emphasize that behind the factorial enhancements of operators is a correlation between operator’s derivative numbers and perturbative orders. Indeed, for the operator $O_{n}$ with $n-1$ $\partial^{2}$ inside, it is the bubble-chain diagrams with bubble numbers around $n$ that causes the factorial growth of $O_{n}$ in $n$. For bubble numbers $n+k\gg n$, the UV region is suppressed by $\frac{1}{k^{n}}$ and will not cause factorial growth. The $O_{n}$ themselves have no Borel ambiguities when regarded as perturbative expansions in $\frac{g}{m^{2}}$.
    The situation is different from marginal theory: as least based on lessons from large-$N$ expansion David (1982, 1984, 1986); Beneke et al. (1998); Marino and Miravitllas (2024); Liu (2024), in asymptotically free theories there are ambiguities to define high-dimensional operators, but these ambiguities can be resolved easily in an “one-for-all” manner using a Borel prescription. Once the Borel prescription has been fixed, there are no further factorial enhancement of “high-twist” operators.
@@
    We should also mention that unlike the renormalons in supper-renormalizable theories, the factorial enhancements observed here are rather subtle and is generated from $\frac{1}{\epsilon}\times\epsilon$ effects. It would be interesting to see if there are simple explanations for the observed factorial growths.
@@
    Finally, we must emphasize that the factorial enhancement is only $n!$. This is sufficient to make the momentum space OPE divergent. However, in the coordinate space expansion there are always additional $\frac{1}{(n!)^{2}}$ suppression factor for the $(z^{2})^{n}$ power, thus a single $n!$ can not affect the convergence of coordinate space OPE or power expansion. Moreover, beyond the $\frac{1}{N}$ expansion, there can be factorial growth in the numbers of Feynman-diagram. But as far as the numbers of Feynman diagrams at order $g^{n}$ are bounded by $n!\times C^{n}$, the factorial enhancement for the $n$-the power in the coefficient functions would at worst lead to $\frac{n!\times n!}{(n!)^{2}}\tilde{C}^{n}(gz^{2})^{n\pm 1}$, which still has a finite convergence radius. This is consistent with the statements in Lukyanov and Zamolodchikov (1997).
@@ Appendix A Expansion of the sunrise diagram
From the above, it is clear that all the singularities have been factorized to the overall factor $(\beta_{1}\beta_{2})^{-s}$, while the rest of the integrand are never singular. Moreover, by partial integrate, ${\cal M}(s)$ becomes a meromorphic function in $s$ with poles at $s=n\in Z$. Furthermore, the ${\cal M}(s)$ decays exponentially as ${\cal O}(e^{-\pi|{\rm Im}(s)|})$ in the vertical strip, while grows in ${\rm Re}(s)$ generically no faster than ${\cal O}\left(9^{{\rm Re}(s)}\right)$. This is consistent with the general pattern of Mellin transform for momentum space correlation function in fixed order perturbation theory.
@@
${\cal M}(s)\bigg{|}_{s\rightarrow 3}\rightarrow\frac{3}{2(s-3)}-\frac{156}{(s-3)^{2}}+\frac{90}{(s-3)^{3}}+\left(60\zeta_{3}+\frac{183}{2}\right)\ .$ (173)

### v2

@@ 4 Comments
    First, we emphasize that due to the off-diagonal cancellation pattern across different powers, the naive power-expansion in $\frac{1}{(p^{2})^{n+2}}$ for the two point function, although free from any ambiguity at any given power, failed to converge. The OPE expansion is also not convergent, unless one redefines all the operators by subtracting out all the factorial growths and adding back to coefficient functions. Even after this procedure, the coefficient function still can not be summed as a power expansion.
@@
    We should emphasize that behind the factorial enhancements of operators is a correlation between operator’s derivative numbers and perturbative orders. Indeed, for the operator $O_{n}$ with $n-1$ $\partial^{2}$ inside, it is the bubble-chain diagrams with bubble numbers around $n$ that cause the factorial growth of $O_{n}$ in $n$: for these diagrams, there is a very large power of the logarithms, while a relatively small algebraic power. For bubble numbers $n+k\gg n$, the UV region is suppressed by $\frac{1}{k^{n}}$ and will no-longer cause factorial growth when $k\sim{\cal O}(n)$. This is consistent with the fact that $O_{n}$ themselves have no Borel ambiguities when regarded as perturbative expansions in $\frac{g}{m^{2}}$.
    The situation is different from marginal theory: as least based on lessons from large-$N$ expansion David (1982, 1984, 1986); Beneke et al. (1998); Marino and Miravitllas (2024); Liu (2024), in asymptotically free theories there are ambiguities to define high-dimensional operators, but these ambiguities can be resolved easily in an “one-for-all” manner using a Borel prescription. Once the Borel prescription has been fixed, there are no further factorial enhancements of “high-twist” operators.
@@
    We should also mention that unlike the renormalons in supper-renormalizable theories, the factorial enhancements observed here are rather subtle and are generated from $\frac{1}{\epsilon}\times\epsilon$ effects. As such, it is less transparent to see its presence using naive $d=2$ running-coupling-type arguments. To see if it works, let’s consider the IR renormalized mass-less bubble chain with
    $F(k^{2})=\frac{1}{4\pi k^{2}}\ln\frac{k^{2}}{\mu^{2}}\ .$ (179)
    Plug it in the diagram with $n$ bubbles, one obtains
    $\frac{-g^{n+1}}{(4\pi)^{n}}\int\frac{d^{d}k}{(2\pi)^{d}}\frac{\mu_{0}^{2-d}}{(k^{2})^{n}}\left(\frac{\mu^{2}}{k^{2}}\right)^{s}\frac{n!}{(p-k)^{2}}\bigg{|}_{s^{n}}-\frac{g^{n+1}}{(4\pi)^{n+1}\epsilon}\frac{1}{(p^{2})^{n}}\ln^{n}\frac{\mu^{2}}{p^{2}}\ .$ (180)
    Here $|_{s^{n}}$ denotes the $s^{n}$ power Taylor expansion coefficient at $s=0$. Notice the last term corresponds to the IR subtraction on the $p-k$ line. In this manner, one obtains
    $-\frac{g^{n+1}n!}{(4\pi)^{n+1}(p^{2})^{n}}\left(\frac{\mu^{2}}{p^{2}}\right)^{s}\frac{\exp(\gamma_{E}\epsilon)\Gamma(-\epsilon)\Gamma(-\epsilon-n-s+1)\Gamma(\epsilon+n+s)}{\Gamma(n+s)\Gamma(-2\epsilon-n-s+1)}\bigg{|}_{s^{n}}$
    $-\frac{g^{n+1}}{(4\pi)^{n+1}\epsilon}\frac{1}{(p^{2})^{n}}\ln^{n}\frac{\mu^{2}}{p^{2}}$
    $\rightarrow\frac{g^{n+1}}{(4\pi)^{n+1}(p^{2})^{n}}\left(\frac{\mu^{2}}{p^{2}}\right)^{s}n!\left(-l_{\mu}+\psi(-n-s+1)+\psi(n+s)+2\gamma_{E}\right)\bigg{|}_{s^{n}}\ .$ (181)
    Again, the $\ln p$ in the $\frac{1}{\epsilon}$ is removed by the IR subtraction on $p-k$, but one is left with a pole at $s=0$. This pole is related to the operator renormalization for $O_{n}$ that contains no dependency on $p$ and can not be calculated accurately in this method. Neglecting this pole, one obtains exactly the Eq. (127) for $C_{n,0}^{B}$ that contains the factorial growth in $n$. Therefore, similar to the 4D bubble chain in QCD Beneke (1999), the naive bubble chain approximation here also produces the correct large $n$ factorial behavior, up to a scheme dependent constant part that is related to operator renormalization and contains no factorial growth in $\overline{\rm MS}$-like schemes.
    The factorial growth in fact allows a simpler explanation: if one consider the integrand
    $\frac{(-1)^{n+1}g^{n+1}}{(4\pi)^{n}}\int_{k^{2}\ll\mu^{2}}\frac{d^{2}k}{(2\pi)^{2}}\frac{1}{(k^{2})^{n}}\ln^{n}\frac{k^{2}}{\mu^{2}}\frac{1}{(p-k)^{2}}\ ,$ (182)
    then for small-$k$, although it has a high degree of IR divergence, the large-power of $k^{2}$ seems to have balanced out the large power of the logarithm and no-factorial behavior seems possible. However, if one Taylor expand the integrand to the first power where the integral converges in the IR, then one has
    $\frac{(-1)^{n+1}g^{n+1}}{(4\pi)^{n}}\int_{k^{2}\ll\mu^{2}}\frac{d^{2}k}{(2\pi)^{2}}\frac{1}{(k^{2})^{n}}\ln^{n}\frac{k^{2}}{\mu^{2}}\frac{(k^{2})^{n}}{(p^{2})^{n+1}}\rightarrow-\frac{g^{n+1}}{(4\pi)^{n}}\frac{\mu^{2}}{p^{2}}\frac{n!}{(p^{2})^{n}}\ ,$ (183)
    which is just the leading non-alternating factorial growth in Eq. (128) after multiplying the overall factor $\frac{g^{n+1}}{(p^{2})^{n}}$. From the above, it is clear that the factorial growth in the coefficient function is produced in the region where $k^{2}\ll\mu^{2}$ in the renormalized bubbles. Naively, in the IR subtraction, $\mu$ serves as the separation point between UV and IR and coefficient functions should not go the momentum region below than $\mu$. The point is, here the subtraction is local in nature (they are just the UV poles for operators) and can not see the momenta within the massless bubble. Even when the $\mu$ is actually larger than the $k^{2}$ of the bubble, one still subtracts up to the point $\mu$ which is actually much harder than the massless bubble. This is completely consistent with the interpretation of the observed factorial growth as an IR renormalon.
@@
    Finally, we must emphasize that the factorial enhancement is only $n!$. This is sufficient to make the momentum space OPE divergent. However, in the coordinate space expansion, there is always an additional $\frac{1}{(n!)^{2}}$ suppression factor for the $(z^{2})^{n}$ power, thus a single $n!$ can not affect the convergence of coordinate space OPE or power expansion at large $N$. Moreover, beyond the $\frac{1}{N}$ expansion, there can be factorial growth in the numbers of Feynman-diagrams. But as far as the number of Feynman diagrams at order $g^{n}$ is bounded by $n!\times C^{n}$, the factorial enhancement for the $n$-the power in the coefficient functions would at worst lead to $\frac{n!\times n!}{(n!)^{2}}\tilde{C}^{n}(gz^{2})^{n\pm 1}$, which still has a finite convergence radius. This is consistent with the statements in Lukyanov and Zamolodchikov (1997).
@@ Appendix A Expansion of the sunrise diagram
From the above, it is clear that all the singularities have been factorized to the overall factor $(\beta_{1}\beta_{2})^{-s}$, while the rest of the integrand are never singular. Moreover, by partial integrating, ${\cal M}(s)$ becomes a meromorphic function in $s$ with poles at $s=n\in Z$. Furthermore, the ${\cal M}(s)$ decays exponentially as ${\cal O}(e^{-\pi|{\rm Im}(s)|})$ in the vertical strip, while grows in ${\rm Re}(s)$ generically no faster than ${\cal O}\left(9^{{\rm Re}(s)}\right)$. This is consistent with the general pattern of Mellin transform for momentum space correlation function in fixed order perturbation theory.
@@
${\cal M}(s)\bigg{|}_{s\rightarrow 3}\rightarrow\frac{3}{2(s-3)}-\frac{156}{(s-3)^{2}}-\frac{90}{(s-3)^{3}}+\left(60\zeta_{3}+\frac{183}{2}\right)\ .$ (199)
@@
Appendix B Angular averages
In this appendix we derive the angular averages for the small-$k$ expansion of the tree-level propagator $\frac{1}{(p-k)^{2}+m^{2}}$ in an arbitrary $d$-dimension.
The simplest way to perform this expansion is to use the following trick. Consider
$I(z,k)=\int\frac{d^{D}p}{(2\pi)^{D}}\frac{e^{ip\cdot z}}{(p-k)^{2}+m^{2}}=\int\frac{d^{D}p}{(2\pi)^{D}}\frac{e^{ip\cdot z}}{p^{2}+m^{2}}e^{ik\cdot z}\ .$ (215)
Clearly, in order to obtain the small $k$ expansion from the region where $p$ is hard, it is sufficient to keep only the hard contribution for the coordinate-space two-point function and then Fourier-transform back. For this purpose, one can write
$\int\frac{d^{D}p}{(2\pi)^{D}}\frac{e^{ip\cdot z}}{p^{2}+m^{2}}\bigg{|}_{\rm hard}=\sum_{l=0}^{\infty}\int\frac{d^{D}p}{(2\pi)^{D}}\frac{(-1)^{l}m^{2l}e^{ip\cdot z}}{(p^{2})^{l+1}}$
$=\sum_{l=0}^{\infty}(-1)^{l}m^{2l}\frac{\Gamma\left(\frac{D}{2}-l-1\right)}{(4\pi)^{\frac{D}{2}}\Gamma(l+1)}\left(\frac{z^{2}}{4}\right)^{1+l-\frac{d}{2}}\ .$ (216)
Now, for the angular average of $e^{ik\cdot z}$, one has
[…]
Given the above, by Fourier transforming back from the coordinate space to the momentum space again, one obtains after angular average over $k$
$\frac{1}{(p-k)^{2}+m^{2}}=\sum_{n=0}^{\infty}\frac{k^{2n}}{(p^{2})^{n+1}}{\cal C}_{n}\left(\frac{m^{2}}{p^{2}}\right)\ ,$ (218)
${\cal C}_{n}\left(\frac{m^{2}}{p^{2}}\right)=\sum_{l=0}^{\infty}\left(\frac{m^{2}}{p^{2}}\right)^{l}\frac{\Gamma\left(1-\epsilon\right)(-1)^{l+n}\Gamma\left(-l-\epsilon\right)(l+n)!}{l!n!\Gamma\left(1-\epsilon+n\right)\Gamma\left(-l-n-\epsilon\right)}$
$=\sum_{l=0}^{\infty}\left(\frac{m^{2}}{p^{2}}\right)^{l}(-1)^{l}\frac{(l+n)!}{l!n!}\frac{\Gamma(1-\epsilon)}{\Gamma(l+1+\epsilon)}\frac{\Gamma(n+l+1+\epsilon)}{\Gamma(n+1-\epsilon)}$
$=\frac{\Gamma(1-\epsilon)\Gamma(n+1+\epsilon)}{\Gamma(1+\epsilon)\Gamma(n+1-\epsilon)}\,_{2}F_{1}\left(n+1,n+1+\epsilon;1+\epsilon;-\frac{m^{2}}{p^{2}}\right)\ .$ (219)
In the large $n$ limit, $C_{n}$ is absent of factorial growth. For example, at $p^{2}=m^{2}$ and $\epsilon=0$, one has
${\cal C}_{n}\bigg{|}_{p^{2}=m^{2},D=2}=\frac{\Gamma\left(\frac{n}{2}+\frac{3}{2}\right)}{\Gamma(n+2)\Gamma\left(-\frac{1}{2}-\frac{n}{2}\right)}\ .$ (220)
It vanishes at $n=2k+1$, while for $n=2k$ it decays exponentially as ${\cal O}\left(2^{-n}\right)$.